FEL Upgrade Project Weekly Brief
 February 5-9, 2001


This week the FEL was operated for continuing the spectroscopic work of G. Luepke for his investigations of the
H defects in Si.  The team was able to improve the wavelength and pointing stability of the FEL to improve his signal-to-noise a factor 4-5 compared to his experimental run in November.


This week F. Dylla and C. Leemann had the pleasure of meeting the Chief of Naval Research, Admiral Jay Cohen at the Feb. 5-6 meeting of the Virginia Research and Technology Commission in Richmond.  During off line discussions Adm. Cohen expressed in his interest in the FEL program and his desire to visit the lab in the near future  for a tour and briefing.

We submitted this week, jointly with VCU, a $3.75 M proposal to the Commonwealth Technology and Research Fund for recommissioning the x-ray synchrotron donated by one of the FEL industrial partners.
Also three proposals went to NSF this week by FEL users to support FEL user lab infrastructure.

WBS 3 (Beam Physics):

B. Yunn evolved an injector solution based on the 500 keV Engwall  gun, sporting very low emittance (~2;3.5 mm-mrad/16;32 pi deg-keV) at high charge state (~135;270 pC). This is adequate for the UV driver and
likely will support JERBIL (D. Douglas's concept for a 4th generation light source).  Work continued on interfacing the  UV bypass, with most effort focussed on management of momentum compaction while avoiding impressively strong trim quads.

WBS 4 (Injector):

Continued gun chamber design for the Upgrade gun. Started planning for gun testing an the end of the 2001.
Gun HVPS - Resolved the differences with the vendor concerning the multiplier stacks.  There is just no other alternative to 2 multiplier stacks and 1 monitor stack.  This means the HVPS tank will have to be
extended for the additional multiplier stack.

WBS 5 (SRF):

 Two (2) cavities fabricated this month; this makes four (4) complete (less groups).
 First article helium vessels are undergoing inspection and QA.
 HOM component design completed and components  are being fabricated.
 Stress test of the beamline flange completed this week.  Flange was deflected 50 mrad and remained
 leak tight at 2K

WBS 6 (RF):

Zone 3 - No progress.

Zone 4 - The system was brought into RF ON for cavity 1 this week.  Work continues on the test software to complete and
document the zone.

Injector RF - Ordered 2 autotransformers as backups for the existing 3 phase Variacs.  PSC is still considering the cost and schedule impact of using autotransformers in the upgrade.  Reliability and space requirements should improve.

WBS 8 (I&C):

The optical BPMs for the upstream optical table, can 2, can 6 and End-of-the-Line Dump are experiencing some problems that are elaborated on in the Optics section of this brief.  The picomotor cable in the  drive laser clean room was also re-terminated to improve reliability.  Two new optical BPM detector assemblies were built and
tested and are available as well as new cables for testing the BPM system.  A test  cable for the new “quad-detectors” has been fabricated, as has a “break-out” cable to monitor and log RS232 “communications”.  Two new local shutter control boxes for user labs are complete and await testing. Also in progress is the construction of power supply racks for each of the user labs. Drawings are in to EECAD for GPIB System drawing and an update to the LSS User Interface Box in the user labs.
WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research continues to model the Small Injector dipole.  They are working the on the 3 D model
    as well as the 2D details.
 Optical Chicane Dipole (DW)
o  We checked the drawing package and started incorporating the information.  We started on the
    specifications for making the magnet.
Arc Dipoles (GY, GX, GQ)
o  AES started the layout the coil's leads, incorporating the 4-in-hand winding and started re directing
    the leads so the water connections could be made while incorporating the bus bars and lead stabilization
o  AES modeled the magnetics of the GY using axisymmetric 2D model which makes provision for the
    out-of-plane curvature of the magnet.  They found that the field is concentrated in the just the inner return leg
    such that the flatness of the field warps to several times the allowed value within the nominal good field width.
    They will now work on a model with increased width of return leg only to see if the flatness returns.
    Their next line of action is to increase the thickness of the pole plates.
o We worked out a way for the leads of the Path Length Correctors to be mounted like the leads of the
    IR Demo's injection dipoles.
QX (3.125" Quad)
o Measurement Probe: The vendor is performing the final machining of the coil support plate coil slots.
o The procurement of steel was started.
o At a working session, we settled on a method of machining fiducial surfaces on the quadrupole that satisfies
    the ability to Wire Electric Discharge Machine the pole tips, provides a mount for the magnets on the test
    stand, allows simple direct mounting to single and double girders and provides for use of a bubble level and
    tooling balls during final alignment.  Sometimes committees actually work!
o Design Package: In parallel with his GW design efforts, N. Vaughan is working on finalizing the QX drawings:
   The drawing package is being modified to incorporate the decisions of
    the working session.
o Prototype Measurement: Effective length measurements were made using a hall probe.
o Manufacturing philosophy: A meeting with Will Oren, head of engineering technicians and shops resulted in a
    decision to have separate procurement of coils and a separate procurement of cores.  Then, once assembly
    drawings are available, we will probably procure assembly from the resulting vendors but keep open the
    option of assembling in-house.
Sextupole with H/V Correctors
o Robin Wines got her magnetic model of the Sextupole to work.  Interpreting data and optimization, if
    necessary, are to follow.

WBS 10 (Wiggler):

Coils for the dispersion section were completed but snowed-in up in New Hampshire.  Coils will be shipped to PECO the end of this week where the core is waiting.  A model for the wiggler is being developed in Radia to
determine the best solution for the field bias issue. Design is progressing on the measurement stand.  No decision has yet been made (Karn and Benson) on whether the existing Group-3 Hall Probes are acceptable or whether a dual axis probe needs to be purchased.

WBS 11 (Optics):
We are finishing review of the check prints for the deformable mirror assembly, and plan to release the drawing package next week.  The design team continues to work on the upgrade optical transport system (OTS) and
optical cavity assemblies.

During operations this week, with extended operation at modest laser power (~ 100 W cw), we found that the O-BPM ADC electronics' limited dynamic range made their use problematic.  We are going to work with I&C
on this, and in the short term install a larger range of neutral density filters.

As mentioned in the Jan 8-12 weekly brief, we replaced the black-anodized Al mirror holders on the collimator linear stage with unanodized holders.  This was done because it appeared that most of the beam drift we observed came from the collimator, and the black-anodized mirror holders, with their high CTE, were a likely candidate.  Our operational experience this week has shown a greatly reduced drift, compared with our run in November.  So, this lends credence to our specifications in the upgrade OTS that all mounts be either temperature stabilized or made of low CTE materials.


The machine availability this week for the continuation of  the user run was good. See the highlights and Optic sections for details.

Privacy and Security Notice